OpenStudy (anonymous):

Joe Thea and Taylor invested in a partnership in the ratio 1:4:7. The partnership is worth 1.6 million and Thea decides to sell her partnership to Joe. How much would njoe need to pay Thea to buy her share?

OpenStudy (ghazi):

joe has to pay 0.53 million dollars x+4x+7x= 1.6 million dollars 12x=1.6 x= 0.133 so share of thea = 4x= 4*0.133= 0.533 million dollars....and joe has to pay this much

OpenStudy (anonymous):

thank you!! :)

OpenStudy (ghazi):

:)